Help please,

I am unsure how to calulate this problem

The following information is available for the Gold Company:

Net income for the year ended December 31, 2014        $127.4

Retained earnings, December 31, 2014                           150.0

Retained earnings, December 31, 2013                           180.0

Total assets, December 31, 2014                                    470.0

Total assets, December 31, 2013                                    442.0

Total liabilities, December 31, 2014                                240.0

Total liabilities, December 31, 2013                                182.0

What is the return on stockholders' equity for the year ended December 31, 2014?

A) 24.1%

B) 27.1%

C) 52.0%

D) 55.4%
Answer #1

Correct answer will be 52 %

Detail working for your refrence

Return on Stockholder Equity= Net Income/ Average Stoccckholder equity
Stockholder Equity= Total Asset- Total Liabilities
2013=442-182=260
2014=470-240=230
Average Equity= (260+230)/2=245
Stock holder Equity=127.40/245= 52%